Posts

Showing posts from May, 2025

NPP Projection Report 5/17 Admin

To do this evening/ was completed: Just a reminder since I don't think I mentioned it in a blog: someone from the Quad Cinema is coming to pick up the Bitter Tears of Petra von Kant print towards the end of May. Lillian or I will update you when we know exactly which day. Print is by the stairs in booth 1 at the moment. Presentation Issues: Preshow Clips/Playlist: First MI8 shows scheduled. I scheduled the content to transfer tomorrow morning at 9:15. Trailer Updates: Special Events/ AV set up: The Women is on house reels. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Printed labels for Barbershop and Abominable Dr. Phibes and attached them to the boxes. If Phibes is able to be packed up Monday night, please schedule a pickup for both on Tuesday otherwise we can send em out on Wednesday. Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): MI8 for...

NPP Projection Close 5/16

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built show playlists for 5/23 for  Mission: Impossible and the remaining R rated first runs in addition to The Red Shoes and The Colors Within (this last film may need additional material added to the pre-show since there are no trailers prior to the feature) Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Posters for Together , Gabby's Dollhouse , and The Bad Guys 2  were delivered and stored behind 3/7 Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked all cues for the weekend and cross referenced the schedule both days to be certain everything was in its right place + initialed for confirmation  Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional...

NPP Projection Admin 5/16

To do this evening/ was completed: New preshow content downloaded and ingested. Most A1 playlist built but one of the R-rated trailer packs needs the Fix Your Heart trailer ingested. Mission Impossible feature ingested. Tech'd Cleopatra Jones and Dr Phibes. Volumes logged. Phibes reel 1 has a bit of a prominent crackle on the soundtrack but it goes away after a bit and all of the other reels are unaffected. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Colors Within, MI8 ingested Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 5/15

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The Howling event in theater 1 went very smoothly (the host ultimately did not need a mic to amplify his acoustic guitar and stood with the mic stand while he played before the Q&A) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): We have yet to receive any feature content for  Mission: Impossible so please be on the lookout for it on the Deluxe portal in the coming days Outgoing/Shipping (content): Print/video Inspection: Re-inspected  Barbershop in 4/6 and placed all reels back on plastic and in the box they were shipped in Got word from a friend at Quad that their projectionist will be emailing us tomorrow about picking up the Petra von Kant print in person, so no labels will be forwarded or UPS return will need to be scheduled Asset tracking doc updated/checked?: Maintenance/cleaning: A guest marked seat 1423 with some so...

NPP Projection Report 5/15 Admin

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Tech for  Cleopatra Jones  for Shouting at the Screen is on the calendar tomorrow, remember to update the playlist. This has been filled in on the cue sheet for next Thursday, but has not yet been looked at on screen. Downloaded and ingested MI8 and Lilo and Stitch preshows. I'm sure we will need to download more when trailer lists get sent out by Kat tomorrow. Lilo playlist built and scheduled for next Thurs, and that's been filled in on the cue sheet. Trailer Updates: Removed Bring Her Back trailer from all playlists which had it and replaced with 28 Years Later . We're not going to open Bring Her Back . Special Events/ AV set up: There will be a musical element to the intro for The Howling tonight. Someone will be playing an acoustic guitar. They will be using mics as usual, all they will require in addition to the two mics is a mic stand. Ordering/Receiving (posters & trailers): Ord...

NPP Projection Close 5/14

To do this evening/ was completed: Film Feast went smoothly. Spotlight remote returned to its container. Barbershop also went smoothly. Mic/mixer broke down afterward. Same for trees mic/cable. Was not able to get to the 35mm letters tonight due to the rain and running around for the second half of the evening. Built Totoro playlist with the english dub DCP (EN-EN). Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 5/14 Admin

To do this evening/ was completed: We had a seat left out in T7 last night. Please remember to put back the ADA seats once the screening is over. Thanks! - L Staff screenings have been emailed out, built/scheduled on the LMS and blocked on the calendar. - L Presentation Issues: Preshow Clips/Playlist: Updated the playlist for and tech'd Paris is Burning . The 1.37/1.33 masking cue was set to 1.66 so I set it back to its usual place. - L R-rated playlists swapped to 5/16 versions for tomorrow, so we're not showing FD6 and Friendship trailers after they've already opened. -A Built Abominable Dr. Phibes and scheduled it. -A Trailer Updates: Special Events/ AV set up: Private event went well in T6 this morning Content removed from the server.  - L Mic set in T4 for the intro tonight. There is also an intro in T3 for the Film Feast but they won't use a mic. Be sure you switch on the spotlight switch for the appropriate theater before the intro and remember to switch it o...

NPP Projection 5/13 Close

To do this evening/ was completed: Presentation Issues: The final round in theater 2 did not start as scheduled and needed to be played manually + no guests arrived for the final round in theater 5 and this show was ended early Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up:  Petra von Kant screening in theater 1 this evening went very well + mics were returned to charge and both projectors were cleaned following the show (vents are closed) Ordering/Receiving (posters & trailers): Eddington posters were delivered and stored behind 3/7 Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Re-inspected  Petra von Kant  after the screening and scheduled it for a return pickup tomorrow morning Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked all cues for tomorrow, cross referenced the schedule, and initialed for confirmation (second round in theater 5 needed...

NPP Projection Report 5/13 Admin

To do this evening/ was completed: Cue sheets made for the coming week with a few exceptions ( Mission Impossible).  Ran Batman and Robin for Desmond. Loops ran afterward for Petra. Looked at The Menu slide and feature, updated the feature volume. Scanners ingested to LMS. Presentation Issues: Preshow Clips/Playlist: Built and scheduled the playlist for the private event in 6 tomorrow morning -A Trailer Updates: Special Events/ AV set up: Mufasa private event went smoothly for the most part. I got here early to check the feature on screen, as it was decided yesterday that we were best renting and downloading from Amazon rather than screening the DCP we made of the DVD. Everything was well and good until I went to start the movie for the audience, and the damn play controls wouldn't go away. I ended up having to quit the app, restart it, and play the movie again. Everything was fine after that. -A Cleaned the projectors for tonight's show. -A Ordering/Receiving (posters & t...

NPP Projection Close 5/12

To do this evening/ was completed: Scheduled the coming week in LMS with a few exceptions (content not available yet). Eraserhead went very smoothly. Applause at the end. Print will be re-inspected and packaged up tomorrow. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 5/12 (Lillian & Will)

To do this evening/ was completed: Will has notes below on Mufasa but I've let Alec know that we'll be streaming this for tomorrow. I downloaded the Prime Video App in T1 for Windows. The film has been rented and downloaded so that there's no chance of it cutting out or behaving strangely. All good for tomorrow. - Lillian  Presentation Issues: Preshow Clips/Playlist: Petra playlist and a few others added to/adjusted on the LMS. Hubba Hubba preshow transferred. - Lillian  Trailer Updates: Special Events/ AV set up: Ran house checks (not an A/V thing but the masking in theater 1 is suffering from a bit of squeakiness) -WB Ordering/Receiving (posters & trailers): Ordering/Receiving (content):  Waiting on the keys for Final D6 and Friendship. Keys for The Howling will open up tomorrow!  - L Outgoing/Shipping (content): Print/video Inspection: Tech'd Mufasa for the private event in theater 1 this morning...it certainly looks like we ripped a standard definition DV...

NPP Projection Open 5/11

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled a show playlist for  The Menu  on 5/14 (which will likely need to be updated and tech'd once John Woods finalizes the event memo) Added the  Barbershop  pre-show to the schedule + placed the feature for The Howling and its corresponding volume/lights cues to its show playlist and scheduled it Scheduled show playlists for Final Destination: Bloodlines  and Friendship  on 5/15 ( Final Destination   credits timing might be off by a few minutes on the cues) Trailer Updates: Special Events/ AV set up: Hairspray and News from Home  concluded their screenings for the weekend in theaters 2 and 4 Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Mufasa DCP finished converting this afternoon and is saved to the Headquarters hard drive which is back in 4/6 (had some difficulty getting the projector in booth 1 to recognize this so pleas...

NPP Projection Report 5/10 Open

To do this evening/ was completed: Regarding hard drive troubles: I spent the better part of the day today working on this, so I'll try to break it all down as clearly as possible. -For starters, we can write to the NTFS drives again. -Most of our external hard drives have always been NTFS formatted. This is the only hard drive format which is read by both the LMS and IMSes. LMS will read exFAT drives as well but that's it. -NTFS is natively Windows-only, but we use a third-party app, called NTFS for Mac, to allow us to read and write to NTFS hard drives on Mac computers (both the 4/6 and T1 iMacs have this installed). -The iOS update must have reset permissions for third-party apps, as NTFS for Mac had some necessary permissions disabled which I believe kept causing the app to crash every time I tried to open it. -I saw the app's icon on the top bar of the screen (which indicates it's running in the background), but every time I tried to actually open the application i...

NPP Projection Close 5/9

To do this evening/ was completed:  Presentation Issues: Preshow Clips/Playlist: Built all rated R first run show playlists for the week of 5/16 and the Paris is Burning  special event playlist for brunch on 5/17-18 (this will need to be tech'd for credits/lights cues) There was not enough space remaining on the George Ivan drive to successfully copy all the "SE/Brunch PreShow" content to it from the P&T drive (that folder requires some 500 extra GB and George Ivan has a little over 200 to spare) so I refrained from partially copying it and/or having the content split on a separate drive and left it as is Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Removed Until Dawn and The Shrouds  and added letters for Bonjour Tristess...

Projection Admin 5/9

To do this evening/ was completed: The rain has been persistent, so we'll be late on changing the marquee and posters. Hopefully it'll subside by tomorrow morning.  This morning Jack was having issues with the computer in 4/6 as all external drives were readable but not writeable. All of these drives were formatted as NTFS and I've been trying to figure out if the OS update has anything to do with compatibility for reading/writing, because they should kind of work in that state on a mac (but some evidence online suggests otherwise and it says Windows NTFS so, you know, that's not ideal) and they want you to use APFS for Ventura (which is what the computer is on now) but exFAT is what we want to use the drives on both macs and PCs. We didn't need anything off the George Ivan drive so I reformatted it for exFAT which made it readable/writeable. The other two external drives, preshows/trailers and VHS Vault, are still NTFS. I copied over the Alice Jr. DCP to George Iva...

NPP Projection Admin 5/8

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Having Fun with...  screening and Q&A in theater 1 went very smoothly + the afterparty got started at Trees and John Woods was on site to assist throughout the evening (note: the Q&A was recorded and saved to the booth 1 desktop and can be extracted/emailed to John Woods whenever) The hosts of the  Tremors screening in 7 requested that we push back the start of the pre-show by 15 minutes and the playlist was moved accordingly to accommodate them + Rebecca helped run the event while I was busy in booth 1 and the film played without any issues Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded and ingested new keys for The Menu Outgoing/Shipping (content): Print/video Inspection: Ran the DVD copy of Mufasa  through MakeMKV and Handbrake...audio is currently being analyzed in DCP-o-matic and this can be convert...

Projection Admin 5/8

To do this evening/ was completed: Our OS in 4/6 has been updated and the Signiant App (for media shuttle deliveries) should be working now. Things look a tiny bit different, really just a small visual shift with the graphics, but that's about it. We're hoping to partner with Signiant/Media Shuttle for future film festival needs (in short - they make the DCPs and we receive/download them). Lots of teching tomorrow. I added a staff screening of News From Home in T4, where we will tech at 1pm. The show will end 13 minutes before we're to tech.  Rebecca (MOD) has the 4k BluRay of  Tremors and will make sure it gets back to the host. Mind you, we have another version in the booth but it is ours.  Presentation Issues: Preshow Clips/Playlist: Built the playlists for Final Destination 6 and Friendship .  I scheduled Bonjour Tristesse on the LMS, built out the formula on the cue sheets and filled in the times for the week of 5/9.  FYI- not sure why I didn't receive...

NPP Projection Close 5/7

To do this evening/ was completed: Blues Run The Game screening and Q&A went smoothly. Mics returned to the booth. The webcam would not cooperate with me, however. I turned it on and off multiple times and could never get an image on screen. Went to device settings and restarted the camera application, got a notification saying that the webcam had to be given access in the privacy settings of the camera app, which I then disabled and re-enabled to no effect, so. Inspected the rest of The Women. Excellent condition. It is a B&W film, but reel three features a technicolor segment. Don't be alarmed, it's part of the film. Put up new 35mm letters as well as the text John sent over for the Doc and Roll screening tomorrow night. He included a colon in the text but I don't think we have that marquee tile. I left all of the brunch side letters in 3/7 booth.  Bonjour Tristesse staff playlist shows up red on the scheduler, saying the feature file is corrupt. I put the film on...

Projection Admin 5/7

To do this evening/ was completed: Hey y'all I saw this morning that there was another slide made yesterday for Having Fun with A-holes, but I already made one on Monday so that it would be ready for the tech on Tues. I did put this info in a thread but added it to the blog. Just be sure to double check when you read through the day's work so that nothing gets repeated. Thanks! - L Staff Screenings have been emailed out, blocked on the calendar and built out on the LMS. For tonight, feel free to add the formula for Bonjour T to the cue sheets. We'll have to schedule the open hours shows on the LMS. - Lillian  Presentation Issues: Preshow Clips/Playlist: Built the playlists: Bitter Tears, Barbershop, Bonjour Tristesse - Lillian   Trailer Updates: Special Events/ AV set up: All SE/private events are ready to go this evening: bring the mics down for the Blues Run the Game screening in T2. John will help with chairs.  I Killed My Mother is good to go in T4.  Private S...

NPP Projection Close 5/6

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Added/transferred the show playlist for Blues Run the Game to theater 2 for its screening tomorrow evening Still waiting on pre-show and feature file elements for Bonjour Tristesse  and Friendship  so these show playlists can be built and scheduled in the coming days Trailer Updates: Special Events/ AV set up: Set up for trivia at Trees Lounge Dance, Girl, Dance intro and screening in theater 4 this evening went very smoothly Ordering/Receiving (posters & trailers): Received a new poster for  Thunderbolts and/or  The New Avengers Ordering/Receiving (content): Received slides for the upcoming screenings of Barbershop and  The Menu  and converted them to DCP/ingested them to the LMS Began converting Cleopatra Jones  to DCP in booth 1 and saved it to the Headquarters drive Outgoing/Shipping (content): Print/video Inspection: Re-inspected  Dance, Girl, Dance after the ...

NPP Projection Admin 5/6

To do this evening/ was completed: Re-inspected Good Night and Good Luck, packaged it up and scheduled an Express pickup for tomorrow 11 - 6 pm. Ran a reel of Eraserhead for Cristina, volume logged.  Inspected the first two reels of The Women. Excellent condition so far. Tech'd Blues Run The Game, volume logged, playlist built. Playlist also built for HFWA. Friendship feature ingested to LMS, A1 playlist still needs to be built and scheduled. New 35mm letters up on the marquee. Booth 5 vacuumed. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ran loops for tonight and cleaned the projectors. Mic tested. -A Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Cleopatra Jones is converting on Handbrake - MC Outgoing/Shipping (content): Print/video Inspection: Ted M. came by to look at his film for Thursday ( Having Fun with A-holes ). He was satisfied with how it looked and sounded. -A Asset tracking doc updated/checked?:...

NPP Projection Close 5/5

To do this evening/ was completed: Scheduled the coming week in LMS with the exception of the first runs we've yet to receive. Good Night and Good Luck went smoothly. Print will be re-inspected tomorrow. Robocop also went fine. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 5/5

To do this evening/ was completed: Hey everyone - the gate broke today, so Vlad has been directing folks outside to come in through the side door where a manager will be waiting to greet them. The maintenance person drove out from Long Island and is working on it. No update yet. - Lil As for our AC Units, I checked the 3, 7 and 5, 5 being the only AC unit that was on and without airflow. Apparently (and this makes sense, but) if the thermostat reads the room as the correct temperature its been set to, it won't blow, but I've just never seen one of the units paused, so I wanted to double check. Brought the temp down in T5 just a hair, and checked to make sure air flow was steady. - Lil Presentation Issues: Preshow Clips/Playlist: Transferred Hairspray, I Killed My Mother, Tremors, News From Home to their respective screening rooms.  Looked at all of these (just not Hairspray as it's encrypted and the others are not) and updated info in the playlists for all. Built the play...

NPP Projection Close 5/4

To do this evening/ was completed: Batman and Robin stills uploaded to Drive. Downloaded and ingested the HubbaHubba and Lynch preshow elements. Playlists built for News From Home and Eraserhead. Barbershop on house reels, I Killed My Mother transferred to T4. Round 2 T7 tomorrow is a manual start but I put the playlist on the LMS schedule anyway, just FYI.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:
To do this evening/ was completed: Presentation Issues: Minor crisis averted earlier today as Thunderbolts OCAP was not scheduled on Screenwriter/none of its pre-show or feature content was present on the theater 7 server. Thankfully everything transferred faster than expected and I was able to run the playlist after skipping ahead in the pre-show. Ultimately the guests missed about ten minutes of it but I alerted the MODs who had orders placed as the remaining content finished transferring Preshow Clips/Playlist: Scheduled the show playlist for Robocop in theater 2 for tomorrow evening's screening and began transferring its feature content to the server Built and scheduled a pre-show playlist for Good Night and Good Luck in theater 1 (still waiting on its Broadway trailer and slide but have put the NH logo slide in its place for the intro in the meantime) and filled out the rest of the cues for this Double checked the show playlist for  Robocop for tomorrow evening and added a lig...

Projection Close 5/4

To do this evening/ was completed: Inspected the print of The Bitter Tears of Petra Von Kant. We'll tech this sometime this week, I reckon. In good condition.  I've added a bunch of photos to the drive of this and Eraser Head.  No issues this evening. Quiet night on my end.  Presentation Issues: Preshow Clips/Playlist: Still waiting on the Broadway Trailer/Slide for Good Night & Good Luck. Reached out to Desmond since this screens on Monday.  Continued SE playlists for this week and the next. Ingested all NoBudge DCPs to the LMS.  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 5/3 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: I made DCPs of the slides for the Bitter Tears intro. There's 3 DCPs, two with the "host" slide (one which fades in and one which fades out), and one with the "cue" slide (no fades) which should go in between the two "host" slide DCPs: 1. Slide_Intro_1_BitterTears_FadeIn 2. Slide_Cue_BitterTears 3. Slide_Intro_2_BitterTears_FadeOut Trailer Updates: Special Events/ AV set up: Postcards went nicely today. Put the mic away after. Print is packed up, I emailed Sony to schedule a pickup and attached the report. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): New download of Alice Junior worked, it's ingested to the LMS. Ingested I Killed My Mother and Tremors DCPs, too. Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Mic input 2 is back in the wall in Trees. Tested it again this mo...

NPP Projection Close 5/2

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Added a feature volume cue to the pre-show playlist for Postcards from the Edge Trailer Updates: Special Events/ AV set up: Finished converting all NoBudge45 shorts to DCP Ordering/Receiving (posters & trailers): Fantastic 4 posters were delivered in addition to teaser posters for The Smashing Machine + a box containing two Lilo & Stitch standees was dropped off and can likely be given away at trivia/to whoever wants a life size Stitch in their home and/or eventually thrown away Ordering/Receiving (content): The Bitter Tears of Petra Von Kant print was delivered and brought up to booth 1 this afternoon + downloaded both slides for this film's pre-show and saved them to the 4/6 desktop File elements for the  Alice Junior DCP are downloading to Jack's webpage and should finish later this evening Outgoing/Shipping (content): Print/video Inspection: Double checked the grease cue marks on reels 3...

NPP Projection Admin 5/2

To do this evening/ was completed: Looked at SoF print and Dance Girl Dance with Cristina. Volumes logged. The SoF print A Day Without Policemen does indeed feature a sort of flowing light fade, at least on the first reel. You can see the color go in and out throughout the reel, not prominently. ALSO: The first reel jumps out of frame a few times, starting with right after the AGFA logo at the beginning, so you'll have to keep an eye on that. T3 programming screening went smoothly. Old forecasts and A1 playlists deleted. New forecasts and preshow content ingested, 5/9 A1 playlists built. Gave T1 a vacuum beneath the curtain, got some detritus off the ground. Tremors is still transcoding in T1 and will be done later this evening. I Killed My Mother has been transferred to the same drive and both can be ingested once Tremors is finished. Looked at Fame in T3. Lights up cue added, and I updated the volume to 6.3. I think this is the right level, as the dialogue mix is pretty quiet an...

Projection Close 5/1

To do this evening/ was completed: Spoons went great, but please know that you should leave the wireless mic mixer off until the show begins. Everything can remain off until it's showtime. Cleaned everything up after. Super neat under there, just needs to be vacuumed in the morning, which I'll do tomorrow.  Presentation Issues: Preshow Clips/Playlist:  Trailer Updates: Special Events/ AV set up: Began DCPing the NoBudge shorts in 4/6.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Received the keys for Robocop .  Downloaded the rest of the NoBudge shorts this evening and put them on a separate drive so we can DCP whilst T1 is being used.  I Killed My Mother finished transcoding this evening. I then began the DCP for Tremors, which I rekcon should finish up tomorrow.  Outgoing/Shipping (content): Print/video Inspection: Inspected the last three reels of Eraserhead .  Asset tracking doc updated/checked?: Maintenance/cleaning...

NPP Projection Admin 5/1 (Alec & Will)

To do this evening/ was completed: Please make the minimum effort of coiling the mic cables after events. The mic in T3 from In the Loop last night was just placed on top of the speaker with the cable falling on the floor, peeking out from under the curtain. This is a great way to make sure these cables get tangled together over time and create lots of unnecessary frustration for people trying to use them in the future. -A Presentation Issues: Preshow Clips/Playlist: Converted the intro slide for  Dance, Girl, Dance to DCP and uploaded it to the LMS + built the pre-show for this screening and transferred it to theater 4 -WB I switched the R rated 4/25 A1 playlists for the 5/2 playlists so we're not showing the Thunderbolts trailer, which opens today. -A Built and scheduled the playlist for the Sinners private event next week. -A Trailer Updates: Special Events/ AV set up: Set up and tested all gear for STB in theater 1 this morning. -WB I noticed the Having Fun With A-Holes DCP...